# SSTP: Efficient Sample Selection for Trajectory Prediction

### v2

@@ 4.3 Ablation Study
Figure 2: Top: Visualization examples of different driven scenarios in trajectory prediction dataset, using Argoverse 1 as an example. Bottom: Distribution of scene density before (left) and after (right) applying SSTP on Argoverse 1 with 50% retention ratio.
Data Retention Ratio. To examine the impact of different data retention ratios on model performance, we conducted experiments with retention rates of $\alpha$ = {60, 50, 40, 30, 20, 10}as shown in Figure 3. When higher model performance is required, retaining 50% of the data already achieves results comparable to training with on full dataset. Specifically, minADE and minFDE increase only slightly from 0.69 to 0.70, from 1.03 to 1.06, while the MR metric remains at a low level with 0.11. This demonstrates the effectiveness of our SSTP method, as the selected 50% subset is of higher quality compared to equally sized subsets selected by other methods. Furthermore, under limited computational resources, retaining only 20% of the data still yields reasonably good results. Although minADE and minFDE increase compared to the 50% subset (from 0.7063 to 0.7491 and from 1.0734 to 1.1666, respectively), the MR metric remains at 0.13. This indicates that even when the dataset size is reduced to $\frac{1}{5}$, our method can still maintain a reasonable level of predictive performance.
@@
4.4 Efficiency
Our method significantly reduces computational time while maintaining strong performance, as shown in Figure 1. On a single NVIDIA RTX 4090, training the HiVT model on the full dataset requires 7.78 hours, achieving a minADE of 0.69. In contrast, utilizing our SSTP method to select a 50% subset requires only 2.17 hours, significantly reducing the overall training time. When training on the selected subset, the total training time decreases to 6.37 hours (2.17 + 4.20), with only a minor increase of 0.01 in minADE. For the HPNet model, full dataset training takes 307.2 hours, whereas training with the selected subset reduces the training time by over 100 hours. This decisively confirms the superior efficiency of our approach in balancing training cost and model performance.
Figure 3: Performance of HiVT-64 and HPNet when training with subset chosen by different selection methods. Comparing at varying data retention ratios with minADE and minFDE. Methods include Random, K-means Clustering, Herding and SSTP, with the full dataset serving as a reference baseline. Lower values indicated better performance.
Backbone Ratio(%) HiVT-64 HPNet
minADE minFDE xMRx minADE minFDE xMRx
HiVT-64 50 0.704 1.073 0.111 0.661 0.913 0.074
60 0.703 1.065 0.111 0.654 0.901 0.072
HPNet 50 0.708 1.079 0.111 0.664 0.917 0.074
60 0.703 1.067 0.108 0.657 0.910 0.075
Table 7: Performance comparison of different backbone models, HiVT-64 and HPNet, trained on selected data at varying retention rates (50% and 60%). The table reports minADE, minFDE, and MR for both models under different data selection strategies.
4.5 Generalizability Study
Different Backbones. To further assess the generalizability of our method, we evaluated its performance using different backbone models. Specifically, we employed pre-trained HiVT-64 and HPNet as feature extractors on the Argoverse 1 dataset and conducted experiments at various data retention ratios. In Table 7, we report the results for the 50% and 60% retention ratios. Additional results for other retention ratios can be found in the Supplementary Material. As shown, regardless of the backbone used for subset selection, the final trajectory prediction performance remained nearly identical. These results suggest that our subset selection strategy is largely independent of the specific feature extractor, highlighting its robustness. This broad applicability makes it a promising approach for optimizing trajectory prediction models across diverse architectures.
Partition minADE minFDE MR
@@ 4.5 Generalizability Study
Table 8: Comparison of data selected by SSTP with different partition intervals $\tau$ and its impact on HiVT performance using the Argoverse 1 dataset.
5 Conclusion
Limitations. Despite the effectiveness of our proposed method, improvements could be made to further reduce the computation overhead of our sample selection process. Further optimization can be made to streamline this process. Additionally, investigating methods to sustain model performance under extremely low data retention rates, e.g., 10%, presents a promising path for future research.

### v3

@@ A.3 Ablation Study
[…]
@@
Table 12: Comparison of SSTP with different partition intervals $\tau$ and its impact on HiVT performance using the Argoverse 1.
Figure 6: Visualization of different agent density scenarios in Argoverse 1.
Figure 7: Visualization of different agent density scenarios in Argoverse 2.
A.4 Qualitative Results
In Figure 6 and Figure 7, we present trajectory samples from three different scene densities in Argoverse 1 and Argoverse 2. In the top row, the scenes feature only a few agents, making them relatively straightforward for the model to predict. Because low-density scenarios dominate the dataset, the model tends to be biased toward these simpler cases, and the lower interaction complexity naturally reduces motion-forecasting uncertainty. In the middle rows, moderate-density scenarios display an increased number of agents. Here, interactions among vehicles become more frequent, creating added complexity for trajectory prediction. In the bottom row, high-density scenes pose significant challenges for trajectory prediction models. These crowded urban intersections and multi-agent interactions are underrepresented in standard datasets. The large number of dynamic agents in these scenarios compounds uncertainty, making it harder for the model to produce accurate predictions. Yet, these are precisely the scenarios that are most critical for ensuring safe autonomous driving. Our SSTP method addresses this imbalance head-on by emphasizing high-density samples. Through a more balanced yet compact selection of training data, the model becomes well-prepared for both common and complex scenarios. As a result, SSTP bolsters model robustness in high-density environments, leading to more reliable trajectory predictions in real-world urban traffic conditions.
LLM Usage
We used ChatGPT solely for grammar correction and LaTeX formatting. They were not involved in research ideation, experiment design, or data analysis. All scientific contributions, methodology, and results are entirely the work of the authors.